# Inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities

### v1

@@ 6 Appendix
Consider $N\geq 2$, $b\in(0,N)$ and $q\geq 1$. If $N=2$ assume also that $q\geq b$. Then $C^{\infty}_{c}(\mathbb{R}^{N})$ is dense in $E^{q}_{b}$.
@@
Assume now $N=2$ and $b\in(0,2)$. If $\max\{b,1\}\leq q\leq 2$, since $|T_{k}(s)|\leq\min\{|s|,k\}$ and $u\in L^{q}_{b}(\mathbb{R}^{2})$, we have
@@
Thus, for $N=2$, $b\in(0,2)$ and $q\geq\max\{b,1\}$ or $N\geq 3$, $b\in(0,N)$ and $q\geq 1$, from (6.1) we obtain $\nabla u_{k}\to\nabla u$ in $(L^{2}(\mathbb{R}^{N}))^{N}$. Therefore, $u_{k}\to u$ in $E^{q}_{b}$.
@@
Since $b\in(0,N)$ we have $w(x):=|x|^{-b}$ a function in $L^{1}_{loc}(\mathbb{R}^{N})$. So, $w\chi_{\{w\leq n^{q}\}}\to w$ in $L^{1}(B_{2k_{0}})$ as $n\to\infty$ and we fix $n_{0}\in\mathbb{N}$ such that

### v2

@@ 6 Appendix
Consider $N\geq 2$, $b\in[0,N)$ and $q\geq 1$. If $N=2$ assume also that $q\geq b$. Then $C^{\infty}_{c}(\mathbb{R}^{N})$ is dense in $E^{q}_{b}$.
@@
Assume now $N=2$ and $b\in[0,2)$. If $\max\{b,1\}\leq q\leq 2$, since $|T_{k}(s)|\leq\min\{|s|,k\}$ and $u\in L^{q}_{b}(\mathbb{R}^{2})$, we have
@@
Thus, for $N=2$, $b\in[0,2)$ and $q\geq\max\{b,1\}$ or $N\geq 3$, $b\in[0,N)$ and $q\geq 1$, from (6.1) we obtain $\nabla u_{k}\to\nabla u$ in $(L^{2}(\mathbb{R}^{N}))^{N}$. Therefore, $u_{k}\to u$ in $E^{q}_{b}$.
@@
Since $b\in[0,N)$ we have $w(x):=|x|^{-b}$ a function in $L^{1}_{loc}(\mathbb{R}^{N})$. So, $w\chi_{\{w\leq n^{q}\}}\to w$ in $L^{1}(B_{2k_{0}})$ as $n\to\infty$ and we fix $n_{0}\in\mathbb{N}$ such that